OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL!!! solve: 8x - 2(x + 1) = 7x + 8 A. 6 B. -6 C. 10 D. -10

OpenStudy (anonymous):

8x-2x-2=7x+8 +2 +2 6x=7x+10 -x=10 x=-10 ITS D